Given the function g(x)=(x+7^3

a) Find the critical points of g(x)

b) On what intervals is g(x) increasing and decreasing

c) At what points, if any does g(x) local and absolute minimum and maximum values?


1
Expert's answer
2021-06-15T18:19:31-0400
"g(x)=(x+7)^3"

Domain: "(-\infin, \infin)"

1.

Find the first derivative with respect to "x"


"g'(x)=((x+7)^3)'=3(x+7)^2"

Find the critical number(s)


"g'(x)=0=>3(x+7)^2=0=>x=-7."

Critical number: -7.

If "x<-7, g'(x)>0, g(x)" increases.


If "x>-7, g'(x)>0, g(x)" increases.


The critical point "x=-7."


b)

The function "g(x)" increases on "(-\infin, \infin)."

The function "g(x)" is never decreasing.


c)

The function "g(x)" is always increasing on "(-\infin, \infin)."

The function "g(x)" has neither maximum nor minimum at "x=-7."

The function "g(x)" has no absolute maxium value on "(-\infin, \infin)."

The function "g(x)" has no absolute minium value on "(-\infin, \infin)."